LOG-IN
Displaying reports 141-160 of 3360.Go to page Start 4 5 6 7 8 9 10 11 12 End
R&D (Speed meter)
Print this report.
YoheiNishino - 11:55, Tuesday 15 October 2024 (3783)Get code to link to this report
Open loop gain measurement of the GR cavity loop

Nishino,

This is a result on 2024/10/10.

I measured the open loop transfer function (OLTF, Fig.1) of the GR lock loop. The filter funtion and actuator efficiency is measured independently (Fig. 2 and 3, Fig. 4 is their sum).

Setting of SR560 (Fig. 7) and Moku (Fig. 8):

SR560: 3 Hz, first order low pass, gain=100

Moku PID: P=10 dB, D=4.118 kHz, DS=13.0 dB

The inferred optical gain is plotted in Fig. 5. From this result, I modified the shape of filter by Moku PID controller as:

P = 10 dB, D = 20 dB, DS = 8 kHz.

Then the modified OLTF is plotted in Fig. 6 with UGF=4 kHz and phase margin=50 degrees.

Images attached to this report
3783_20241015044504_oltf.png 3783_20241015044514_lpfsasrmoku.png 3783_20241015044523_pztamp.png 3783_20241015044528_lpfsasrmokupztamp.png 3783_20241015044533_h.png 3783_20241015044609_oltfmodified.png 3783_20241015045223_img0237.jpeg 3783_20241015045229_img0238.png
General (General)
Print this report.
ShalikaSingh - 10:24, Wednesday 09 October 2024 (3777)Get code to link to this report
Old elogs are looking strange

It seems there is some issue with all elogs made before the server change. They are looking weird. For instance https://gw-elog.mtk.nao.ac.jp/osl/index.php?r=3342

Due to this issue, sometimes the search filter fails, because it is unable to find the word(since it looks strange). The issue happens only if I search and go to the elog. Perhaps not if I go directly without searching through keyword option. 

Images attached to this report
3777_20241009032547_screenshot20241009102640.png
BIGFOOT (Characterization)
Print this report.
ShalikaSingh - 16:21, Tuesday 08 October 2024 (3773)Get code to link to this report
Comment to Re-measuring BS birefringence (Click here to view original report: 3766)

I made a mistake in measuring the BS. I rotated only 90deg. It was supposed to be rotated by 180 deg.

BIGFOOT (cell birefringence)
Print this report.
MarcEisenmann - 10:16, Monday 07 October 2024 (3772)Get code to link to this report
adapter for force gauge

I designed some adapters between force gauge and hydrogel holders.

They're 3D printed with help of Kanzawa-san and models are saved in the dropbox/mount and holder folder.

BIGFOOT (Characterization)
Print this report.
ShalikaSingh - 08:14, Monday 07 October 2024 (3771)Get code to link to this report
Comment to S7 birefringence measurement (Click here to view original report: 3753)

The birefringence obtained is Fig 1. Since there were too few points (112/6050) after filtering. I will re take measurements. 

Images attached to this comment
3771_20241007011455_s7ret.png
BIGFOOT (cell birefringence)
Print this report.
ShalikaSingh - 08:13, Monday 07 October 2024 (3770)Get code to link to this report
Comment to Hydrogel measurement with displacement (Click here to view original report: 3760)

The diattenuation vs. Displacement is as in Fig 1. Doesn't change with increasing displacement. 

Images attached to this comment
3770_20241007011318_diattenuationvsdisplacement.png
BIGFOOT (Characterization)
Print this report.
ShalikaSingh - 08:05, Monday 07 October 2024 (3767)Get code to link to this report
Comment to Re-measuring BS birefringence (Click here to view original report: 3766)

The birefringence of BS changed from 17deg to 10 deg after alignment tuning. See the picture for its Jones matrix (Fig1) and birefringence (Fig2). As you can see the JM doesn't have -ve sign any more in the real parts, and is closer to identiy matrix.

The condition on physical realizablity had error bar of 1E-4 (to lower two peaks on BS birefringence).

Polarization ray tracing to correct coordinate systems seems correct. Perhaps can tune BS alignment a bit more. 

Images attached to this comment
3767_20241007010421_bsjm.png 3767_20241007010426_bsret.png
BIGFOOT (Characterization)
Print this report.
ShalikaSingh - 18:36, Friday 04 October 2024 (3766)Get code to link to this report
Re-measuring BS birefringence

Owing to the spurious 17deg of circular retardation observed in BS birefringence characterisitics after correction using Polarization ray tracing, I redid calibration.

1. First it seems the alignment was off by 0.2 cm before the BS. I corrected this using the last steering mirror before the LC path. The beam height is now 7.5cm all along. The alignment was also checked with BS and was slightly(0.1cm) tuned to obtain good alignment in its reflection. 

2. The BS is removed. LC voltage changes from 0-3.5V with 0.1V step and 100 avg.

Polarization states without BS:

C:\Users\atama\Dropbox\LC-Experiment\Measurement Data\Polarization states\20241004\Fri, Oct 4, 2024 6-35-14 PM.txt

3. The BS is rotated by 180 degrees to measure its back face.

Polarization states after relfection from back side of BS:

C:\Users\atama\Dropbox\LC-Experiment\Measurement Data\Birefringence Measurements\Beam Splitter Calibration\20241006\back_refl\Sun, Oct 6, 2024 5-32-19 PM.txt

Comments related to this report
ShalikaSingh - 08:05, Monday 07 October 2024 (3767)

The birefringence of BS changed from 17deg to 10 deg after alignment tuning. See the picture for its Jones matrix (Fig1) and birefringence (Fig2). As you can see the JM doesn't have -ve sign any more in the real parts, and is closer to identiy matrix.

The condition on physical realizablity had error bar of 1E-4 (to lower two peaks on BS birefringence).

Polarization ray tracing to correct coordinate systems seems correct. Perhaps can tune BS alignment a bit more. 

Images attached to this comment
3767_20241007010421_bsjm.png 3767_20241007010426_bsret.png
ShalikaSingh - 16:21, Tuesday 08 October 2024 (3773)

I made a mistake in measuring the BS. I rotated only 90deg. It was supposed to be rotated by 180 deg.

BIGFOOT (Readout)
Print this report.
ShalikaSingh - 18:00, Tuesday 01 October 2024 (3765)Get code to link to this report
Comment to RF circuit for EOM (Click here to view original report: 3745)

I also added a voltage regulator 7815 (for +15V) and 79 (for -15V) to supply voltage for opamp. A 0.33uF ceramic capacitor is added on the input side of these regulators. Another 0.1uF ceramic capacitor is added on the output side of these capacitors.

BIGFOOT (Readout)
Print this report.
ShalikaSingh - 18:23, Friday 27 September 2024 (3764)Get code to link to this report
Comment to RF circuit for EOM (Click here to view original report: 3745)

I can use a feedback resistance of 1.1k instead of 1.07k, since 1.1k is readily available. Also, the inductors measure about 6.4mH(using multimeter), instead of 6.8mH. In series they will be 12.8mH instead of 13.6m in the deisgn. This shifts the resonant frequency by 18kHz, but doesn't change the gain or the current properties of the circuit too much. So, even if the components will not be perfect we will not overshoot the 20mA limit of the opAmp current at the output.

The OpAmp in elec shop is Op27G / OP27 whose max current is 20mA. 

There is a discrepancy in the reported output current for OP27. Texas instruments says 30mA in their datasheet.  On the other hand Digikey reports 20mA. 

R&D (FilterCavity)
Print this report.
MichaelPage - 18:44, Wednesday 25 September 2024 (3763)Get code to link to this report
Transfer functions and gain setting of SHG/GRMC

From last Friday

I previously attempted to lock the SHG by switching around the DDS1 DAC channels - in the normal configuration DAC0 outputs -9 dBm to +14 dB -> 5 dBm to main laser EOM, while DAC1 outputs 9 dBm to splitter for SHG/IRMC demod. I switched them around but this just makes the SHG oscillate even at 0 gain. So I put them back to normal. The rationale for switching is that the wiki shows an 88 MHz Qubig EOM datasheet which specifies that we need about 12-15 dBm RF power to achieve 0.2 modulation index. However, this is for the OPO EOM "EO-88K3-NIR", while the ML EOM is "PM8-NIR-88". I cannot find the datasheet for this particular EOM but other PM8 EOM datasheets indicate that 5 dBm is sufficient to generate 0.2 modulation index.

I measured the open loop transfer function EPS1/EPS2 of the SHG (fig 1, 2) and GRMC (fig 3, 4) using FFT mode of the spectrum analyzer and injecting about 100 mW noise to PERTURB IN. The coherence is basically 1 around the unity gain frequency. I also checked the SHG transfer function with swept sine mode and it gives basically the same result (fig 5). The SHG unity gain frequency is a bit low. Nominally it should be between 2-3 kHz (Aritomi thesis) but it sits about 1.6 kHz with 50 degrees phase margin when set to 0.09 gain. However, increasing the gain further past 0.1 makes the SHG oscillate quite a lot, even though the sticker on the servo says 0.2 indicating a former optimal setting. I set the GRMC gain to 0.8 which gives 3.4 kHz UGF with 45 degrees phase margin.

The GRMC error signal is clearly visible but also a bit noisy. It's quite curious that the GRMC signal shows up at 88 MHz (fig 6, 7), which is the modulation applied to the IR carrier, despite also propagating through three 1064 reflection dichroics. But evidently it has worked to lock the GRMC since the EOM replacement a while ago. I have a bit of difficulty in locking the GRMC to TEM00 though, a lot of the time it just locks to near zero transmission amplitude and I have to keep hitting the rest button a lot. when I managed to lock it with 4.2 MZ offset (stabilizing GRMC transmission to 25 mW) and noted the GRMC transmission level (GRMC TRANSMISS IN) of 316 mV. I set the GRMC threshold to about 100 mV which helps for avoiding HOMs but it's still not reliable for locking to TEM00.

Images attached to this report
3763_20240925114355_shgoltffft.jpeg 3763_20240925114400_shgoltfcoherence.jpeg 3763_20240925114406_grmcoltf.jpeg 3763_20240925114413_grmcoltfcoherence.jpeg 3763_20240925114422_shgoltfswept.jpeg 3763_20240925114432_grmcreflrf1.jpeg 3763_20240925114436_grmcreflrf2.jpeg
BIGFOOT (cell birefringence)
Print this report.
ShalikaSingh - 18:42, Wednesday 25 September 2024 (3762)Get code to link to this report
Comment to Hydrogel measurement with displacement (Click here to view original report: 3760)

The force constant k was evaluated as follows:

mg = kx

k = (delta mg) / (delta x) = ((0.06-0.03) * 9.8)/(0.01 * (28-27))  = 29.4

BIGFOOT (cell birefringence)
Print this report.
ShalikaSingh - 18:38, Wednesday 25 September 2024 (3761)Get code to link to this report
Comment to Hydrogel measurement with displacement (Click here to view original report: 3760)

The retardation of the data was evaluated and plotted. We can see a linear dependency of retardation(effective and linear retardation) with the stretch of hydrogel. On the other hand, circular retardation and linear retardation along 45deg axis doesn't change.

The distibution of retardation was fit with gaussian fit and the mean and std was estimated. Then for each displacement the analysis was repeated to make the plot 1.

Plot 1 Retardation vs. Displacement

Images attached to this comment
3761_20240925113805_retardationvsdisplacement.png
BIGFOOT (cell birefringence)
Print this report.
ShalikaSingh - 14:19, Wednesday 25 September 2024 (3760)Get code to link to this report
Hydrogel measurement with displacement

[Katsuki, Shalika, Marc]

The thickness of the hydrogel is 1.5mm. The measurement is taken with LC voltage from 0.5V - 3 V with 0.25V step and 10avg.

26.7 cm (F=0g)

C:\Users\atama\Dropbox\LC-Experiment\Measurement Data\cell birefringence\hydrogel\Wed, Sep 25, 2024 10-42-00 AM.txt

26.8 cm

C:\Users\atama\Dropbox\LC-Experiment\Measurement Data\cell birefringence\hydrogel\Wed, Sep 25, 2024 11-45-27 AM.txt

26.9 cm

C:\Users\atama\Dropbox\LC-Experiment\Measurement Data\cell birefringence\hydrogel\Wed, Sep 25, 2024 11-48-42 AM.txt

27 cm

C:\Users\atama\Dropbox\LC-Experiment\Measurement Data\cell birefringence\hydrogel\Wed, Sep 25, 2024 11-52-24 AM.txt

27.1 cm

C:\Users\atama\Dropbox\LC-Experiment\Measurement Data\cell birefringence\hydrogel\Wed, Sep 25, 2024 11-56-57 AM.txt

27.2 cm

C:\Users\atama\Dropbox\LC-Experiment\Measurement Data\cell birefringence\hydrogel\Wed, Sep 25, 2024 12-01-33 PM.txt

27.3 cm

C:\Users\atama\Dropbox\LC-Experiment\Measurement Data\cell birefringence\hydrogel\Wed, Sep 25, 2024 12-06-09 PM.txt

Hydrogel changed here.

27.4 cm (discarded for analysis)

C:\Users\atama\Dropbox\LC-Experiment\Measurement Data\cell birefringence\hydrogel\Wed, Sep 25, 2024 1-46-34 PM.txt

27.5 cm (discarded for analysis)

C:\Users\atama\Dropbox\LC-Experiment\Measurement Data\cell birefringence\hydrogel\Wed, Sep 25, 2024 1-52-02 PM.txt

Comments related to this report
ShalikaSingh - 18:38, Wednesday 25 September 2024 (3761)

The retardation of the data was evaluated and plotted. We can see a linear dependency of retardation(effective and linear retardation) with the stretch of hydrogel. On the other hand, circular retardation and linear retardation along 45deg axis doesn't change.

The distibution of retardation was fit with gaussian fit and the mean and std was estimated. Then for each displacement the analysis was repeated to make the plot 1.

Plot 1 Retardation vs. Displacement

Images attached to this comment
3761_20240925113805_retardationvsdisplacement.png
ShalikaSingh - 18:42, Wednesday 25 September 2024 (3762)

The force constant k was evaluated as follows:

mg = kx

k = (delta mg) / (delta x) = ((0.06-0.03) * 9.8)/(0.01 * (28-27))  = 29.4

ShalikaSingh - 08:13, Monday 07 October 2024 (3770)

The diattenuation vs. Displacement is as in Fig 1. Doesn't change with increasing displacement. 

Images attached to this comment
3770_20241007011318_diattenuationvsdisplacement.png
BIGFOOT (Characterization)
Print this report.
ShalikaSingh - 13:28, Wednesday 25 September 2024 (3753)Get code to link to this report
S7 birefringence measurement

The beam in the reflection of beam splitter was tuned to remove the misalignment of the beam (beam height was chaning by 1cm at the end of the table). Now the beam height is same as before the BS.

I placed the Shinkosha S7 after the beam splitter reflection and before the camera. The LC voltages are changed from 0.5V - 3V with 0.25V step size, with 10 averaging.

without sample

C:\Users\atama\Dropbox\LC-Experiment\Measurement Data\Birefringence Measurements\S7\Tue, Sep 24, 2024 3-23-26 PM.txt

with S7

C:\Users\atama\Dropbox\LC-Experiment\Measurement Data\Birefringence Measurements\S7\Wed, Sep 25, 2024 10-07-47 AM.txt

Comments related to this report
ShalikaSingh - 08:14, Monday 07 October 2024 (3771)

The birefringence obtained is Fig 1. Since there were too few points (112/6050) after filtering. I will re take measurements. 

Images attached to this comment
3771_20241007011455_s7ret.png
BIGFOOT (cell birefringence)
Print this report.
ShalikaSingh - 20:24, Tuesday 24 September 2024 (3757)Get code to link to this report
hydrogel birefringence measurement

[Katsuki, Marc, Shalika]

A 1mm thin hydrogel was prepared between plates to measure its birefringence. We change the LC voltage from 0.5-3 V with step size of 0.25V and averaging  of 10th order.

without sample

C:\Users\atama\Dropbox\LC-Experiment\Measurement Data\cell birefringence\hydrogel\Tue, Sep 24, 2024 3-23-26 PM.txt

We try to shake the hydrogel to see if any movement due to air flow will change the output. The range of azimuth change is 0.02 deg and ell change is 0.1 deg. We then try to stretch the hydrogel using force gauge. The minimum reading of force gauge is 10g.

1. At, F = 10g

C:\Users\atama\Dropbox\LC-Experiment\Measurement Data\cell birefringence\hydrogel\Tue, Sep 24, 2024 3-27-28 PM.txt

It seemed there were issues related to the resolution of the gauge, and we weren't able to see anything other than 10. When we saw the force gauge reading of 40g, the hydrogel broke. We started taking measurements related to the position of the black actuator on the bar

2. 31 cm : C:\Users\atama\Dropbox\LC-Experiment\Measurement Data\cell birefringence\hydrogel\Tue, Sep 24, 2024 4-02-55 PM.txt

3. 30 cm: C:\Users\atama\Dropbox\LC-Experiment\Measurement Data\cell birefringence\hydrogel\Tue, Sep 24, 2024 4-22-05 PM.txt

4. 29.5 cm : C:\Users\atama\Dropbox\LC-Experiment\Measurement Data\cell birefringence\hydrogel\Tue, Sep 24, 2024 4-16-38 PM.txt

 

The figures are in order of the measurement taken.

Images attached to this report
3757_20240924132426_f10g.png 3757_20240924132433_l31cm.png 3757_20240924132438_l30cm.png 3757_20240924132443_5cm.png
BIGFOOT (General)
Print this report.
ShalikaSingh - 17:09, Tuesday 24 September 2024 (3756)Get code to link to this report
Comment to Beam characterization for EOM (Click here to view original report: 3754)

The lens actually was f=50mm and f=-50mm. I have changed the beam characteristics accordingly.

BIGFOOT (General)
Print this report.
ShalikaSingh - 13:26, Tuesday 24 September 2024 (3755)Get code to link to this report
Comment to Beam characterization for EOM (Click here to view original report: 3754)

The lens change was mentioned in 3563

BIGFOOT (General)
Print this report.
ShalikaSingh - 11:58, Tuesday 24 September 2024 (3754)Get code to link to this report
Beam characterization for EOM

There is a scale on the table kept in backward direction, i.e increasing distance from laser is decreasing measurement on the scale. 

In reference to elog 3561 the beam was as attached in Fig 1. The beam characterization was originally done with f=100mm lens at 101.5cm. The waist of beam before the lens was at -47.1 mm, i.e at 101.5 cm +4.71 cm=106.21 cm

According to the elog the lens f = 100 mm was moved later by d=0.2m. But, now the setup has two f=50mm lens. This was never mentioned in elogs that the lens was changed. 

For the modified setup:

The two f = 50 mm, -50mm lens are placed at 96.5 cm and 90.5cm.Hence, the relative distance between lens and waist of beam is 9.71cm and 15.71cm respectively.

The new beam propagation is as in Fig 2. I propagated the original beam (one before f=100 mm lens) after the two f = 50mm lens. 

The beam power before filter is 64mW. 

I wish to replace the energy filter before the HWP on the EOM path, with a beam splitter or beam sampler and make a separate path for SLM. 

Images attached to this report
3754_20240924045757_completebeambeforeandafterf100mm.png 3754_20240924100951_completebeambeforeandafter2f50mm.png
Comments related to this report
ShalikaSingh - 13:26, Tuesday 24 September 2024 (3755)

The lens change was mentioned in 3563

ShalikaSingh - 17:09, Tuesday 24 September 2024 (3756)

The lens actually was f=50mm and f=-50mm. I have changed the beam characteristics accordingly.

BIGFOOT (Readout)
Print this report.
ShalikaSingh - 15:22, Friday 20 September 2024 (3750)Get code to link to this report
Comment to RF circuit for EOM (Click here to view original report: 3745)

The Vpi, half wave voltage for the EOM is defined by the equation

V_{pi} = 0.361 * lambda - 23.844 = 0.361 * 1064 - 23.844 = 360.26 V (considering crystal with 14pF capacitane). The one we have says a capacitance of 12pF. 

(where lambda is in nm)

I tuned the circuit a bit more to obtain the gain of 7.397k with Moku supplying a 0.05V. So, with the updated circuit we will have

1. 0.05*7.397k = 369.8 V

2. Impedance of 48.9 ohm

3. OpAmp current of 11 mA.

4. Resonant frequency of 0.39MHz

I have updated the circuit diagram and transfer function.

PS: Moku can provide minimum of 1mVp-p using waveform generator function. 

Images attached to this comment
3750_20240920082206_check2ckt.png 3750_20240920082211_check2tf.png